# Oversmoothing, “Oversquashing”, Heterophily, Long-Range, and more: Demystifying Common Beliefs in Graph Machine Learning

### v2

@@ Abstract
After a renaissance phase in which researchers revisited the message-passing paradigm through the lens of deep learning, the graph machine learning community shifted its attention towards a deeper and practical understanding of message-passing’s benefits and limitations. In this position paper, we notice how the fast pace of progress around the topics of oversmoothing and oversquashing, the homophily-heterophily dichotomy, and long-range tasks, came with the consolidation of commonly accepted beliefs and assumptions that are not always true nor easy to distinguish from each other. We argue that this has led to ambiguities around the investigated problems, preventing researchers from focusing on and addressing precise research questions while causing a good amount of misunderstandings. Our contribution wants to make such common beliefs explicit and encourage critical thinking around these topics, supported by simple but noteworthy counterexamples. The hope is to clarify the distinction between the different issues and promote separate but intertwined research directions to address them.
@@ 1 Introduction
The last decade has seen an increasing scholarly interest in machine learning for graph-structured data [90, 71, 43, 10]. After an initial focus on the design of various message-passing architectures [38], inheriting from the recurrent [85] and convolutional [70] Deep Graph Networks (DGN), together with the analysis of their expressive power [103, 72], researchers later turned their attention to the intrinsic limitations of the message-passing strategy and the relation between the graph, the task, and the attainable performance. We refer, in particular, to the fact that node embeddings may become increasingly similar to each other as more message-passing layers are used [84], the loss of information that results from aggregating too many messages onto a single node embedding [4], the presence of topological bottlenecks [93], the existence of neighbors of different classes [98], and the propagation of information between far ends of a graph [30]. Addressing these limitations makes a difference when applying message-passing models, including foundational ones [14], to large and topologically varying graphs at different scales, from proteins with hundreds of thousands of atoms [107] to dynamically evolving social networks [62] with billions of users, where such limitations manifest together.
@@
In this position paper, we argue and provide evidence that this is potentially the case for the afore-mentioned issues of oversmoothing (OSM), oversquashing [^1] We discuss and address the troubling abuse of the term “oversquashing” in later sections. (OSQ), heterophily, and long-range dependencies, driving researchers away by the difficulty of circumventing common beliefs while concurrently introducing novel contributions. Scientific progress is therefore slowed down both in terms of reduced workforce and clarity of the problems to be addressed; failure to acknowledge existing inconsistencies may well lead to a reiterated spreading of questionable claims.
@@
While reviewing the literature, we identify and isolate nine common beliefs that, in our opinion, cause great confusion and ambiguities in the field. We then demystify such beliefs by providing simple and possibly memorable counterexamples that should be easy to recall. By encouraging critical thinking around these issues and separating the different research questions, we hope to foster further advancements in the graph machine learning field.
Disclaimer:
@@
We remark that the goal of this position paper is not explicitly pinpointing criticalities in previous works; on the contrary, these works were fundamental to forming our current understanding and ultimately producing this manuscript. We also acknowledge that the list of referenced works cannot be exhaustive, due to the field’s size, and it mainly serves to support our arguments.
@@
Table 1 summarizes our findings about common beliefs in the literature, together with the list of papers where we could find mentions of them [^2] Sometimes we found overly general claims in the first sections, later refined, which nonetheless contribute to the spreading of common beliefs.. We logically divide common beliefs about OSM, OSQ, and the homophily-heterophily dichotomy. In the following sections, we discuss each belief, provide counterexamples, and summarize our arguments with take-home messages.
Common Belief References
OSM 1. OSM is the cause of performance        degradation. [57, 46, 18, 24, 22, 47, 49, 59, 81, 108, 110, 97, 111, 17, 25, 50, 83, 95, 19, 68, 74, 52, 84, 96, 100, 32, 51, 82, 91, 99]
2. OSM is a property of all DGNs [102, 37, 56, 24, 46, 47, 49, 108, 110, 4, 111, 1, 6, 25, 50, 53, 93, 83, 3, 19, 29, 34, 41, 74, 96, 84, 87, 100, 101, 8, 9, 35, 51, 82, 7, 99]
Hom-Het 3. Homophily is good, heterophily is bad. [77, 110, 112, 58, 64, 97, 6, 17, 39, 61, 78, 15, 8, 9, 42, 82, 109]
4. Long-range propagation is evaluated on       heterophilic graphs. [6, 95, 3, 16, 68, 48, 41, 8, 94]
5. Different classes imply different features. [2, 77, 58, 64, 67, 92, 65, 84, 8, 9, 15, 48, 98, 109]
OSQ 6. OSQ synonym of a topological bottleneck. [102, 6, 12, 25, 28, 92, 93, 95, 3, 11, 16, 29, 34, 36, 41, 52, 60, 74, 87, 88, 96, 105, 8, 9, 13, 40, 35, 48, 51, 89, 7, 44, 91, 94]
7. OSQ synonym of computational        bottleneck. [4, 1, 6, 12, 25, 28, 92, 93, 95, 3, 11, 16, 30, 34, 41, 45, 74, 52, 87, 88, 96, 8, 9, 13, 48, 89, 7, 44, 91]
8. OSQ problematic for long-range tasks. [4, 1, 12, 25, 28, 93, 95, 3, 16, 19, 29, 34, 36, 52, 74, 60, 88, 105, 13, 40, 35, 48, 89, 9, 7, 44, 91]
9. Topological bottlenecks associated with        long-range problems. [93, 95, 3, 16, 52, 60, 88, 35]
@@
For readers that are less familiar with the topic and its definitions, Appendix A.1 provides a brief introduction to message-passing models.
2 Is Oversmoothing Really a Problem?
@@
Oversmoothing broadly refers to the phenomenon where, as we stack more message-passing layers in a DGN, the node embeddings become increasingly similar to each other eventually collapsing into a low-dimensional [49, 75]–or even single-vector [18]–subspace. This creates an almost constant representation, independent of the original node-feature distribution, and can potentially result in loss of discriminative power along the way [57, 18, 75].
@@
Formally, et $H^{\ell}\in\mathbb{R}^{n\times d}$ be the matrix of node embeddings after $\ell$ message-passing layers in a DGN, where $n$ is the number of nodes and $d$ is the hidden dimension. Consider a similarity (or separation) function $\pi\colon\mathcal{H}\to\mathbb{R}$, where $\mathcal{H}=\mathbb{R}^{n\times d}$ is the space of all possible embedding matrices. We say that a DGN experiences OSM if
@@
where $c$ is some constant indicating a collapse of embeddings. Deviation metrics (e.g., Dirichlet Energy [18], MAD [22]) and subspace-collapse criteria [75, 48, 82] all measure the same OSM intuition: as depth increases, node embeddings shrink toward a nearly constant subspace.
@@
A widespread claim in the literature is that OSM happens regardless of the specific architecture or the underlying graph. Early theoretical works support this view by analyzing message-passing propagation as a diffusion process: iterated normalized-Laplacian updates converge to a degree-weighted stationary distribution [18, 41], while heat-kernel diffusion converges to a constant vector [75, 5]. The resulting bounds quantify the rate of OSM in terms of the singular values of the feature transform $W$ and the eigenvalues of the graph structure $G$.
These conclusions, however, rely on restrictive assumptions. Later work has relaxed them by introducing learnable feature transforms, non-linear activations, and more elaborate architectures [75, 18, 100], yet no existing proof shows inevitable collapse under realistic training regimes. In practice, remedies such as residual/skip connections, normalization layers, or gating mechanisms are explicitly architectural changes designed to maintain local distinctions, calling into question the universality of this OSM claim.
In addition, many studies probe OSM with untrained (weights frozen at initialization) linear GCN stacks, which is an experimental choice that hides the effect of learning and may lead to the wrong conclusions, as noted by Zhang et al. [106]. Indeed, Cong et al. [27, Figure  2] report OSM only for frozen-weight networks; once parameters are allowed to adapt, the models preserve informative variance.
Together, these observations suggest that OSM is not an inevitable consequence of message-passing but rather a contingent outcome that depends on training dynamics and architectural design choices.
@@
We show a simple training scenario where we see how OSM is not a property of all DGNs and how different elements make it difficult to draw clear conclusions. We train several DGNs under two propagation variants: the vanilla $AXW$ update and the rescaled $AX(2W)$, inspired by Roth and Liebig [82, Figure 1]. We measure OSM with 2 different metrics: Dirichlet Energy (DE) and its norm-normalized version, the Rayleigh Coefficient (RQ), which was also previously used in some works [18, 39, 82, 68]. Figure 1 shows three key facts: (i) some architectures never collapse, (ii) a minor rescaling can reverse the trend, and (iii) DE and RQ often disagree. Hence OSM is neither universal nor straightforward to diagnose.
@@
First, OSM, as measured by Dirichlet Energy (DE), is not universal: GIN’s DE explodes instead of collapsing in the vanilla setting (a), which shows that changing the aggregation-function may lead to an opposite behavior of OSM effect.
@@
Second, a minor scaling in the feature transformation ($2W$ instead of $W$) flips the behavior of several models: curves that decayed in (a) now grow or stabilize, and vice-versa. Similar small tweaks (normalization layers, self-loops, alternative aggregators) can likewise create or remove DE collapse, which has been leveraged by prior work to propose OSM mitigation approaches, such as the ones based on feature normalization [108, 110].
@@
Finally, whether OSM is observed depends on the measure of choice. DE reflects raw smoothness, whereas the RQ normalizes by the feature norm. As a result, the same model can exhibit mutually contradictory trends for different metrics [106]. First, GCN’s DE collapses under the vanilla aggretation (a), explodes with a simple rescaling (b), yet RQ remains essentially flat in both normalised plots (c–d), indicating that GCN embeddings are being rescaled, not necessarily oversmoothed. On the contrary, GAT and SAGE, which had similar behavior as GCN in (a) and (b), decay using RQ (c-d) at a (roughly) linear rate, highlighting how different architectures respond differently to the use of RQ instead of DE. Furthermore, subfigure (d) shows that $AX(2W)$ can stabilize RQ for certain methods, suggesting that normalizations or small parameter adjustments do not affect all models uniformly. Therefore, conclusions about OSM depend heavily on metric and model, an observation that we rarely found in the literature.
@@
In conclusion, OSM is neither inevitable nor uniquely defined: its observation hinges on the architecture, on seemingly innocuous hyper-parameters, and on which stability metric (DE vs. RQ) one adopts. Therefore, a natural question arises: does OSM actually limit the models’ predictive accuracy? We investigate this question in the next section.
@@
Figure 1: (a-b): We depict the evolution, with increasing number of layers, of the $\mathrm{DE}=\mathrm{tr}(X^{T}\Delta X)$ and the feature norm $\|X\|^{2}_{F}$, using $W$ and $2W$ feature transformations for different architectures. (c-d): Evolution of the $\mathrm{RQ}=\mathrm{tr}(X^{T}\Delta X)/\|X\|^{2}_{F}$ for $W$ and $2W$ as before. Experiments run on the Cora dataset for 50 random seeds. A larger version for better visualization is available in Fig. 8.
@@
Part of the literature focuses on the narrative that OSM is the cause of lower test accuracy in DGNs. The hypothesis is that if embeddings collapse to a non-meaningful space, then the separability of the nodes will become challenging and accuracy decreases.
@@
Regarding the first statement, although it is true that if there is total collapse to the same value then the embeddings will not be informative at all, the main problem remains the node embeddings’ separability. As shown in the previous section, some changes in the architectures can avoid OSM, but they might have no impact on the overall accuracy. For instance, multiplying by two the weight matrix leads to a general increase in DE for all architectures, however, the accuracy will remain the same since the embeddings have been simply scaled up and the embeddings’ separability is not affected negatively. In addition, avoiding an embedding collapse does not necessarily lead to an improvement in generalization accuracy. For instance, comparing a GCN with and without bias, both versions show a decrease in performance as the number of layers increases, whereas only GCN shows a collapse in DE [84, Figure  3].
On the other hand, and related to the second statement, embedding collapse will not always lead to a decrease in accuracy. OSM happens faster in some subspaces than in others and this effect will be beneficial if labels are correlated with those subspaces [53]. For instance, if we classify points into two classes, and all nodes of distinct classes collapse into different points, the OSM metric will detect such collapse. However, the separability of the node embeddings will remain possible, illustrating also the limits of wide-spread used OSM metrics with respect to label information.
This intuitive behavior has been identified in the literature as a form of “beneficial” smoothing phase [53, 82, 101]). In this phase, the nodes of each class first collapse into a class-dependent point before the potential second stage, at which point all nodes converge to the same representation. Finally, although overall pairwise distances might shrink in deep layers, within-class distances might contract more than between-class ones, so class separability improves despite the global collapse, as discussed by Cong et al. [27].
In conclusion, low accuracy in DGNs cannot be attributed to OSM alone. The separability of node embeddings plays a major role, where other training problems such as vanishing gradients or over-fitting also arise when using a big number of layers [108, 27, 104, 7, 76].
Message of the Section i) OSM is not a property of all DGNs ii) OSM is not necessarily the cause of performance degradation. The performance is related to node embeddings’ separability, which can be also affected by many other elements, such as vanishing gradients. iii) Therefore, to study the performance of DGNs, it might be better to study how they achieve separability of node embeddings, and how the OSM relates to node separability.

### v3

@@ Abstract
After a renaissance phase in which researchers revisited the message-passing paradigm through the lens of deep learning, the graph machine learning community shifted its attention towards a deeper and practical understanding of message-passing’s benefits and limitations. In this paper, we notice how the fast pace of progress around the topics of oversmoothing and oversquashing, the homophily-heterophily dichotomy, and long-range tasks, came with the consolidation of commonly accepted beliefs and assumptions – under the form of universal statements – that are not always true nor easy to distinguish from each other. We argue that this has led to ambiguities around the investigated problems, preventing researchers from focusing on and addressing precise research questions while causing a good amount of misunderstandings. Our contribution is to make such common beliefs explicit and encourage critical thinking around these topics, refuting universal statements via simple yet formally sufficient counterexamples. The end goal is to clarify conceptual differences, helping researchers address more clearly defined and targeted problems.
@@ 1 Introduction
The last decade has seen an increasing scholarly interest in machine learning for graph-structured data (Sperduti and Starita, 1997; Micheli and Sestito, 2005; Gori et al., 2005; Bacciu et al., 2020). After an initial focus on the design of various message-passing architectures (Gilmer et al., 2017), inheriting from the recurrent (Scarselli et al., 2009) and convolutional (Micheli, 2009) Deep Graph Networks (DGN), together with the analysis of their expressive power (Xu et al., 2019; Morris et al., 2023), researchers later turned their attention to the intrinsic limitations of the message-passing strategy and the relation between the graph, the task, and the attainable performance. We refer, in particular, to the fact that node embeddings may become increasingly similar to each other as more message-passing layers are used (Rusch et al., 2023), the loss of information that results from aggregating too many messages onto a single node embedding (Alon and Yahav, 2021), the presence of topological bottlenecks (Topping et al., 2022), the existence of neighbors of different classes (Wang et al., 2024), and the propagation of information between far ends of a graph (Dwivedi et al., 2022). Addressing these limitations makes a difference when applying message-passing models, including foundational ones (Beaini et al., 2024), to large and topologically varying graphs at different scales, from proteins with hundreds of thousands of atoms (Zhang et al., 2023) to dynamically evolving social networks (Longa et al., 2023) with billions of users, where such limitations manifest together.
@@
In this paper, we argue and provide evidence that this is potentially the case for the afore-mentioned issues of oversmoothing (OSM), oversquashing (OSQ), [^1] We discuss and address the troubling abuse of the term “oversquashing” in later sections. heterophily, and long-range dependencies, driving researchers away by the difficulty of circumventing common beliefs while concurrently introducing novel contributions. Scientific progress is therefore slowed down both in terms of reduced workforce and clarity of the problems to be addressed; failure to acknowledge existing inconsistencies may well lead to a reiterated spreading of questionable claims.
Contribution:
@@
While reviewing the literature, we identified and grouped together nine common beliefs that cause great confusion and ambiguities in the field, especially when taken separately. We then demystify such beliefs by providing simple counterexamples that should be easy to recall. By encouraging critical thinking around these issues and separating the different research questions, we hope to foster further advancements in the graph machine learning field.
@@
Disclaimer: We remark that the goal of this paper is not explicitly pinpointing criticalities in previous works; on the contrary, these works were fundamental to forming our current understanding and ultimately producing this manuscript. We also acknowledge that the list of referenced works cannot be exhaustive, due to the field’s size, and it mainly serves to support our arguments.
@@
Table 1 summarizes our findings about common beliefs in the literature, together with the list of papers where we could find mentions of them. [^2] Sometimes we found overly general claims in the first sections, later refined, which nonetheless contribute to the spreading of common beliefs. We logically divide common beliefs about OSM, OSQ, and the homophily-heterophily dichotomy. In the following sections, we discuss each belief, provide counterexamples, and summarize our arguments with take-home messages.
@@
For readers that are less familiar with the topic and its definitions, Appendix B.1 provides a brief introduction to message-passing models. In addition, Appendix E provides a discussion with alternative perspectives on the issues discussed in this paper. The code we use to run experiments is available at https://github.com/AdrianArnaiz/demystifyingGraphML.
Table 1: List of common beliefs. For a non-exhaustive list of papers that make those claims, please refer to Table 2 in the Appendix.
Beliefs Beliefs
OSM 1. OSM is the cause of performance degradation. OSQ 6. OSQ synonym of a topological bottleneck.
2. OSM is a property of all DGNs. 7. OSQ synonym of computational bottleneck.
Hom-Het 3. Homophily is good, heterophily is bad. 8. OSQ problematic for long-range tasks as well.
4. Long-range propagation is evaluated on heterophilic graphs. 9. Topological bottlenecks associated with long-range problems.
5. Different classes imply different features.
2 Is Oversmoothing a Real Issue?
@@
Oversmoothing broadly refers to the phenomenon where, as we stack more message-passing layers in a DGN, the node embeddings become increasingly similar to each other, eventually collapsing into a low-dimensional (Huang et al., 2020; Oono and Suzuki, 2020)–or even single-vector (Cai and Wang, 2020)–subspace. This creates an almost constant representation, independent of the original node-feature distribution, and can potentially result in loss of discriminative power (Li et al., 2018; Cai and Wang, 2020; Oono and Suzuki, 2020).
@@
Formally, let $H^{\ell}\in\mathbb{R}^{n\times d}$ be the matrix of node embeddings after $\ell$ message-passing layers in a DGN, where $n$ is the number of nodes and $d$ is the hidden dimension. Consider a similarity (or separation) function $\pi\colon\mathcal{H}\to\mathbb{R}$, where $\mathcal{H}=\mathbb{R}^{n\times d}$ is the space of all possible embedding matrices. We say that a DGN experiences OSM if
@@
where $c$ is some constant indicating a collapse of embeddings. Deviation metrics (e.g., Dirichlet Energy (Cai and Wang, 2020), MAD (Chen et al., 2020a)) and subspace-collapse criteria (Oono and Suzuki, 2020; Huang et al., 2024; Roth and Liebig, 2024) all measure the same intuition: as depth increases, node embeddings shrink toward a nearly constant subspace or degree-proportional vectors.
@@
A widespread claim in the literature is that OSM happens regardless of the specific architecture or the underlying graph. Early theoretical works support this view by analyzing message-passing propagation as a diffusion process: iterated normalized-Laplacian updates converge to a degree-weighted stationary distribution (Cai and Wang, 2020; Giraldo et al., 2023), while heat-kernel diffusion converges to a constant vector (Oono and Suzuki, 2020; Arnaiz-Rodriguez and Velingker, 2024). The resulting bounds quantify the rate of OSM in terms of the singular values of the feature transform $W$ and the eigenvalues of the graph structure $G$.
These conclusions, however, rely on restrictive assumptions. Later work has relaxed them by introducing learnable feature transforms, non-linear activations, and more elaborate architectures (Oono and Suzuki, 2020; Cai and Wang, 2020; Wu et al., 2023a), yet no existing proof shows inevitable collapse under realistic training regimes. In practice, remedies such as residual/skip connections, normalization layers, or gating mechanisms are explicitly architectural changes designed to maintain local distinctions, calling into question the universality of this OSM claim.
In addition, many studies probe OSM with untrained (weights frozen at initialization) linear GCN stacks, an experimental choice that hides the effect of learning and may lead to the wrong conclusions, as noted by Zhang et al. (2025). Indeed, Cong et al. (2021, Figure  2) report OSM only for frozen-weight networks; once parameters start adapting, the models preserve informative variance. Together, these observations suggest that OSM is not an inevitable consequence of message-passing but rather a contingent outcome that depends on training dynamics and architectural design choices.
@@
We show a simple training scenario where we see how OSM is not a property of all DGNs and how different elements make it difficult to draw clear conclusions. We used several DGNs under two propagation variants: the vanilla $AXW$ update and the rescaled $AX(2W)$, following the same experimental procedure as Roth and Liebig (2024, Figure 1). We measure OSM with 2 different metrics: Dirichlet Energy (DE) and its norm-normalized version, the Rayleigh Coefficient (RQ), which was also previously used in some works (Cai and Wang, 2020; Di Giovanni et al., 2023b; Roth and Liebig, 2024; Maskey et al., 2023), defined as:
$\mathrm{DE}(\mathbf{H}^{\ell})$ $=\mathrm{Tr}\!\left((\mathbf{H}^{\ell})^{T}\mathbf{L}\mathbf{H}^{\ell}\right)=\frac{1}{2}\sum_{u,v\in\mathcal{E}}\lVert\mathbf{h}_{u}-\mathbf{h}_{v}\rVert_{2}^{2},\qquad\qquad\mathrm{RQ}=\frac{\mathrm{Tr}\!\left((\mathbf{H}^{\ell})^{T}\mathbf{L}\mathbf{H}^{\ell}\right)}{\lVert\mathbf{H}^{\ell}\rVert_{F}^{2}}$ (2)
where $\mathbf{L}$ is the graph Laplacian and $\mathbf{H}^{\ell}$ is the node embedding matrix at layer $\ell$. DE measures the raw smoothness of the embeddings, while RQ normalizes this smoothness by the overall scale of the embeddings, providing a relative measure of how much the embeddings are collapsing compared to their norm. Appendix B.2 provides more details on these metrics and their relation to OSM.
Figure 1 shows three key facts: (i) some architectures never collapse, (ii) a minor rescaling can reverse the trend, and (iii) DE and RQ often disagree. Hence, OSM is neither universal nor straightforward to diagnose.
@@
First, OSM, as measured by DE, is not universal: GIN’s DE explodes instead of collapsing in the vanilla setting (a), which shows that changing the aggregation-function may lead to an opposite behavior of the OSM effect.
@@
Second, a minor scaling in the feature transformation ($2W$ instead of $W$) flips the behavior of several models: curves that decayed in (a) now grow or stabilize, and vice-versa; this behavior is theoretically analyzed in Di Giovanni et al. (2023b). Similar small tweaks (normalization layers, self-loops, alternative aggregators) can likewise create or remove DE collapse, which has been leveraged by prior work to propose OSM mitigation approaches, such as the ones based on feature normalization (Wu et al., 2019; Zhao and Akoglu, 2020; Zhou et al., 2020).
@@
Finally, whether OSM is observed depends on the measure of choice. DE reflects raw smoothness, whereas the RQ normalizes by the feature norm. As a result, the same model can exhibit mutually contradictory trends for different metrics (Zhang et al., 2025). First, GCN’s DE collapses under the vanilla aggregation (a), explodes with a simple rescaling (b), yet RQ remains essentially flat in both normalised plots (c–d), indicating that GCN embeddings are being rescaled, not necessarily oversmoothed. On the contrary, GAT and SAGE, which had similar behavior as GCN in (a) and (b), decay using RQ (c-d) at a (roughly) linear rate, highlighting how different architectures respond differently to the use of RQ instead of DE. Furthermore, subfigure (d) shows that $AX(2W)$ can stabilize RQ for certain methods, suggesting that normalizations or small parameter adjustments do not affect all models uniformly. Therefore, conclusions about OSM depend heavily on metric and model, an observation that we rarely found in the literature.
@@
In conclusion, OSM is neither inevitable nor uniquely defined: its observation hinges on the architecture, on seemingly innocuous hyper-parameters, and on which stability metric (DE vs. RQ) one adopts (Zhang et al., 2022). Therefore, a natural question arises: Does OSM actually limit the models’ predictive accuracy? We investigate this question in the next section.
@@
Figure 1: (a-b): We depict the evolution, with increasing number of layers, of the $\mathrm{DE}$, using $W$ and $2W$ feature transformations for different architectures. (c-d): Evolution of the $\mathrm{RQ}$ for $W$ and $2W$ as before. Experiments run on the Cora dataset for 50 random seeds. A larger version for better visualization is available in Fig. 8.
@@
Part of the literature focuses on the narrative that OSM is the cause of lower test accuracy in DGNs. The hypothesis is that if embeddings collapse to a non-meaningful space, then the separability of the nodes will become challenging, and accuracy decreases.
@@
Regarding the first statement, although it is true that if there is total collapse to the same value then the embeddings will not be informative at all, the main problem remains the node embeddings’ separability. As shown in the previous section, some changes in the architectures can avoid OSM, but they might have no impact on the overall accuracy. For instance, multiplying by two the weight matrix leads to a general increase in DE for all architectures; however, the accuracy will remain the same since the embeddings have been simply scaled up, and the embeddings’ separability is not affected negatively. In addition, avoiding an embedding collapse does not necessarily lead to an improvement in generalization accuracy. For instance, comparing a GCN with and without bias, both versions show a decrease in performance as the number of layers increases, whereas only GCN shows a collapse in DE (Rusch et al., 2023, Figure 3).
On the other hand, and related to the second statement, embedding collapse will not always lead to a decrease in accuracy. OSM happens faster in some subspaces than in others, and this effect will be beneficial if labels are correlated with those subspaces (Keriven, 2022). For instance, if we classify points into two classes, and all nodes of distinct classes collapse into different points, the OSM metric will detect such a collapse. However, the separability of the node embeddings will remain possible, illustrating also the limits of widely used OSM metrics with respect to label information.
This intuitive behavior has been identified in the literature as a form of “beneficial” smoothing phase (Keriven, 2022; Roth and Liebig, 2024; Wu et al., 2023b). In this phase, the nodes of each class first collapse into a class-dependent point before the potential second stage, at which point all nodes converge to the same representation. Finally, although overall pairwise distances might shrink in deep layers, within-class distances might contract more than between-class ones, so class separability improves despite the global collapse, as discussed by Cong et al. (2021).
In conclusion, low accuracy in DGNs cannot be attributed to OSM alone. The separability of node embeddings plays a major role, where other training problems such as vanishing gradients or over-fitting also arise when using a big number of layers (Zhao and Akoglu, 2020; Cong et al., 2021; Yang et al., 2020; Arroyo et al., 2025; Park et al., 2025).
Call to Action i) Do not assume OSM applies to all DGNs ii) Avoid attributing performance drops to OSM by default. The performance is related to node embeddings’ separability, which can also be affected by many other elements, such as vanishing gradients or over-fitting. iii) Shift research towards achieving separability of node embeddings and, optionally, how OSM relates to that.